# Are AI Detectors Good Enough? A Survey on Quality of Datasets With Machine-Generated Texts

### v2

@@ 5 Experiments
From each dataset, we sampled 1000 documents from the test set, balanced between two classes. Regarding baselines, we fine-tuned mdeberta-v3-base for each dataset and evaluated the model, information about training can be found in the Appendix C. To evaluate the quality of baselines, Binoculars and Fast-DetectGPT, we utilised falcon-rw-1b Almazrouei et al. (2023) and gpt-neo-2.7B Black et al. (2021) respectively. It is worth noting that with the last two methods we were only able to measure quality for samples in English.
Our objective was to show that datasets of lower quality have shifts that will be easily recognised by the models ”from the first step”, hence we have not performed any hyperparameter tuning, only one iteration of fine-tuning and testing of the underlying models. In the experiment with topological features we used roberta-base, just as the authors of the original paper. In the experiment with perturbations and shuffling, the multilingual-e5-large encoder was used to build embeddings of texts, which shows high metrics on encoding high-resource languages Wang et al. (2024a).
6 Results
@@
Dataset BERT Binoculars DetectGPT
GPT-2 0.972 0.495 0.412
TweepFake 0.941 0.845 0.864
HC3 0.998 0.931 0.972
Ghost- Buster 0.910 0.683 0.711
MGT- Bench 0.961 0.364 0.447
MAGE 0.835 0.632 0.654
M4 0.987 0.871 0.881
OutFox 0.901 0.692 0.707
SemEval Mono 0.991 0.913 0.924
SemEval Multi 0.994
RuATD 0.765
DAGPap 0.968 0.333 0.562
PAN24 0.826 0.411 0.890
AuTex En 0.941 0.783 0.911
AuTex Es 0.933
IberAuTex 0.964
MGT-1 Mono 0.904 0.665 0.683
MGT-1 Multi 0.934
Table 3: Classification results with different detectors estimated using $F_{1}$-score. BERT here is DeBERTa-based classifier. Binoculars and DetectGPT work only with English texts, thus we could not apply them to datasets with non-English texts.
The results of the comparison of the designed features on the selected datasets are presented in Table 2. Regarding the PHD and TTS score, in previous works it was shown that texts from language models have smaller PHD values than human-written ones, however this result was obtained for GPT-2, GPT-3.5 and OPT models, and this trend could change for more recent language models, that generate more human-resembling texts. If texts of different origin have high $\text{KL}_{\text{TTS}}$, it means that it is easier for a detector to separate such texts. $\text{KL}_{\text{TTS}}$ is also constrained for shorter texts, see Section 7. As for PHD, we hypothesize that generated texts of good quality should have similar PHD with human-written ones. Additionally, we compare the distributions of PHD for all datasets, see Figure 3. Again, the distributions for texts of both origin should be similar, which holds for texts from SemEval, PAN24 and MGT-1.
@@
Finally, in Table 3 we show the results of applying modern classifiers to the chosen test parts. For instance, on the datasets with low values in Table 2, a quality close to 1 can be achieved, which indicates the clear presence of detector bias towards them, or a structural feature that is too obvious for the model. It is not possible to judge the quality of the data only by achieving $F_{1}$ values close to 1, but by combining the values of the two tables, we can estimate which set has better quality data and which has lower quality data.
@@ 7 Discussion
Regarding $\text{KL}_{\text{TTS}}$, on Fig. 2 we show 4 datasets with high value of it. While GhostBuster and PAN24 received such high score due to discrepancy on texts with higher dimensions, MGTBench and DAGPap22 did it due to the difference in distributions themselves. Note also that $\text{KL}_{\text{TTS}}$ may not perform well with very short texts, since the internal method of computing PHD requires sufficiently long texts for stable computation. Therefore, we discard $\text{KL}_{\text{TTS}}$ on RuATD and AuTex23-es and TweepFake, as they do not fit the criteria, see Table 1. On top of that, it has already been shown that texts must be of sufficient length Gritsay et al. (2022) to build reliable detectors.
Analysing the values in the Table 2, we can trace the presence of sufficiently high quality data in the selected datasets. The developed attributes in aggregate are able to reflect the quality of the generated dataset from different perspectives and angles. We propose to utilise these attributes in combination with other statistical tools for evaluating data quality, e.g., Zipf’s law Powers (1998).
Presented statistics can be utilised to estimate the quality of collections and to improve them. Also, datasets that collect machine-generated content may provide utility for the two more general purposes as well. First, high-quality generated data can be utilised to evaluate the quality of the causal model during training, as one of the training objectives to improve model answers and make it more human-like. Secondly, good detectors can help to clean training sets, as large proportion of low-quality generated texts in those sets can result in emerging biases towards incorrect structure and rubbish fragments in the output of the model in the future.
@@ 8 Conclusion
In the current research, we discussed the problem of quality of datasets with AI-generated texts used for testing corresponding detectors. This problem is relevant, as the quality of test data directly influences the quality of widely used detectors. We conducted a review of datasets from competitions and scientific publications on datasets aimed at the detection of AI-generated content and proposed methods to evaluate the quality of datasets containing AI excerpts based on different structural features. We evaluated topological features, robustness to adversarial attacks, and similarity of attention patterns as estimators of data quality. We concluded that all analysed datasets fail in one another of our methods and do not allow to reliably estimate AI detectors. We encourage researchers to propose their own ways for quality assessment, which will allow to create a comprehensive system of evaluation of the detection datasets. Our work aims to contribute to a better understanding of the difference between human and machine text, which will ultimately contribute to preserving the integrity of information in the world.
@@ 9 Limitations
In our work we focused on the task of binary classification, thus suggested methods are not optimal for the task of detection of the hybrid AI-human content. Also, some methods do not work properly on short texts, however, so do the detectors.
@@ Appendix B Evaluation results of competitions
Table 5 shows the winning scores in the competitions reviewed in this paper. In the AuTex and IberAuTex competitions it was forbidden to use additional data to fine-tune the detection algorithms. In the other collections it was allowed, we can notice a high quality near perfect in them. We should note the low value of metrics on the RuATD dataset, which can be explained by the limited number of available high-quality language models in Russian during the competition.

### v3

@@ 5 Experiments
Dataset DeBERTa Binoculars DetectGPT
GPT-2 0.972 0.495 0.412
HC3 0.998 0.931 0.972
GhostBuster 0.910 0.683 0.711
MGTBench 0.961 0.364 0.447
MAGE 0.835 0.632 0.654
M4 0.987 0.871 0.881
OutFox 0.901 0.692 0.707
TweepFake 0.941 0.845 0.864
SemEval24 Mono 0.991 0.913 0.924
SemEval24 Multi 0.994
RuATD 0.765
DAGPap22 0.968 0.333 0.562
PAN24 0.826 0.411 0.890
AuTex23en 0.941 0.783 0.911
AuTex23es 0.933
IberAuTex 0.964
MGT-1 Mono 0.904 0.665 0.683
MGT-1 Multi 0.934
Table 2: Classification results with different detectors estimated using $F_{1}$-score. Binoculars and DetectGPT work only with English texts, thus we could not apply them to datasets with non-English texts.
@@
From each dataset, we sampled 1000 documents from the test set, balanced between two classes. Regarding baselines, we fine-tuned mdeberta-v3-base for each dataset and evaluated the model. Additional information about hyperparameters during training can be found in the Appendix C. To evaluate the quality of Binoculars and Fast-DetectGPT, we utilised falcon-rw-1b (Almazrouei et al. 2023) and gpt-neo-2.7B (Black et al. 2021) respectively. It is worth noting that with the last two methods we were only able to measure quality for samples in English.
Our objective was to show that datasets of lower quality have shifts that will be easily recognised by the models ”from the first step”, hence we have not performed any hyperparameter tuning, only one iteration of fine-tuning and testing of the underlying models. In the experiment with topological features we used roberta-base, just as the authors of the original paper. In the experiment with perturbations and shuffling, the multilingual-e5-large encoder was used to build embeddings of texts, which shows high metrics on encoding high-resource languages (Wang et al. 2024a).
6 Results
The results of the comparison of the designed features in the selected datasets are presented in Table 3. Regarding the PHD and TTS score, in previous works it was shown that texts from language models have smaller PHD values than human-written ones; however, this result was obtained for GPT-2, GPT-3.5 and OPT models, and this trend could change for more recent language models that generate more human-resembling texts. If texts of different origin have high $\text{KL}_{\text{TTS}}$, it means that it is easier for a detector to separate such texts. $\text{KL}_{\text{TTS}}$ is also constrained for shorter texts, see Section 7. As for PHD, we hypothesise that generated texts of good quality should have PHD similar to human-written ones. Additionally, we compare the distributions of PHD for all datasets on Figure 3. Again, the distributions for texts of both origin should be similar, which mainly holds for texts from SemEval, PAN24 and MGT-1.
@@ 6 Results
Finally, in Table 2 we show the results of applying modern detectors to the chosen test datasets. For instance, on the datasets with low values in Table 3, a quality close to 1 can be achieved, indicating the clear presence of detector bias towards them, or a structural feature that is too obvious for the detection model. It is not possible to judge the quality of the data only by achieving $F_{1}$ values close to 1, but by combining the values of the two tables, we can estimate which set has better quality data and which has lower quality data.
@@ 7 Discussion
Regarding $\text{KL}_{\text{TTS}}$, on Fig. 2 we show 4 datasets with the high value of it. While GhostBuster and PAN24 received such a high score due to the discrepancy on texts with higher dimensions, MGTBench and DAGPap22 did it due to the difference in distributions themselves. Note also that $\text{KL}_{\text{TTS}}$ may not perform well with very short texts, since the internal method of computing PHD requires sufficiently long texts for stable computation. Therefore, we discard $\text{KL}_{\text{TTS}}$ on RuATD and AuTex23-es and Tweepfake, as they do not fit the criteria; see Table 1. In addition, it has already been shown that the texts must be of sufficient length (Gritsay, Grabovoy, and Chekhovich 2022) to build reliable detectors.
Analysing the values from Table 3, we can trace the presence of sufficiently high quality data in the selected datasets. The developed attributes in aggregate are able to reflect the quality of the generated dataset from different perspectives and angles. We propose to utilise these attributes in combination with other statistical tools to evaluate data quality, for example, Zipf’s law (Powers 1998).
Presented statistics can be utilised to estimate the quality of collections and to improve them. In addition, datasets that collect machine-generated content may also provide utility for the two more general purposes. First, high-quality generated data can be utilised to evaluate the quality of the causal model during training, as one of the training objectives to improve model answers and make it more human-like. Secondly, good detectors can help to clean training sets, as large proportion of low-quality generated texts in those sets can result in emerging biases towards incorrect structure and rubbish fragments in the output of the model in the future.
The question of whether poor performance by detectors implies poor dataset quality yields an ambiguous answer. For instance, in (Hans et al. 2024), the Binoculars method achieves an F1-score close to 1.0, while our experiments produced a wide range of scores: from 0.33 on DAGPap22 to 0.93 on HC3. For HC3, all three detectors performed similarly, suggesting that the HC3 texts are relatively easy to detect. However, this consistency does not extend to DAGPap22. For instance, the DeBERTa-based detector achieved an F1-score of 0.96, while DetectGPT scored only 0.562. This pattern, where the DeBERTa-based detector achieves notably higher scores than the other two methods, was observed across a significant portion of the analysed datasets. We attribute this strong performance to the fine-tuning of the DeBERTa-based detector.
Conversely, the low scores for Binoculars merit further scrutiny. Even when focusing on domains specifically tested by its authors, such as PAN24 (News) and Outfox (Student Essays), the scores fall well below the near-perfect results reported in (Hans et al. 2024). This discrepancy suggests that the Binoculars detector may not be representative. Similarly, in our experiments, DetectGPT’s scores are comparable to Binoculars’ scores, potentially indicating similar underlying issues with the robustness of these detectors.
@@ 8 Conclusion
In the current research, we discussed the problem of quality of datasets with AI-generated texts used for testing corresponding detectors. This problem is relevant, as the quality of test data directly influences the quality of widely used detectors. We conducted a review of datasets from competitions and scientific publications on datasets aimed at the detection of AI-generated content and proposed methods to evaluate the quality of datasets containing AI excerpts based on different structural features. We evaluated topological features, robustness to adversarial attacks, and performance of the widely used detectors on these datasets. We concluded that all analysed datasets fail in one or another of our methods and do not allow to reliably estimate AI detectors. We encourage researchers to propose their own ways for quality assessment, which will allow to create a comprehensive system of evaluation of the detection datasets. Our work aims to contribute to a better understanding of the difference between human and machine text, which will ultimately contribute to preserving the integrity of information in the world.
@@ 9 Limitations
In our work we focused on the task of binary classification, thus suggested methods are not optimal for the task of detection of the hybrid AI-human content. Also, some methods do not work properly on short texts, however, this is a known issue for short texts.
@@ Appendix B Evaluation results of competitions
Table 5 shows the winning scores in the competitions reviewed in this paper. In the AuTex and IberAuTex competitions it was forbidden to use additional data to fine-tune the detection algorithms. In the other collections it was allowed, we can notice a high quality near perfect in them. We should note the low value of the metrics on the RuATD dataset, which can be explained by the limited number of high-quality language models available in Russian during the competition.